# Linear and nonlinear instability of vortex columns

### v2

@@ 1.3. Nonlinear instability
coupled with the one-dimensional reduced equation $(I-Q)G=0$, that is,
$\int_{\mathbb{R}}\left(\widetilde{X}_{\rm in}K_{\rm err}(W+\Psi)-R\phi_{\rm out}X_{\rm out}^{\prime\prime}-2n^{-3/4}R\phi_{\rm out}^{\prime}X_{\rm out}^{\prime}\right)W\,\mathrm{d}\xi=0\,.$ (2.19)
@@
$\int\Psi\overline{W}=0\,,$ (2.20)
for example, and $K_{\rm err}$ is small. This dictates some constraints on $\ell_{\rm in},\ell_{\rm out}$ (in addition to (2.2)), see the discussion below Lemma 2.11 for details.
Thus far, we have neglected the outer equation. It will be extended to $\mathbb{R}^{+}$:
@@
and, after we analyze $k$ in the region $|r-r_{0}|\gtrsim n^{-3/4}$, where $\widetilde{k}$ denotes an appropriate extension of $k$ guaranteeing coercivity.
To complete the construction, the projected equation (2.18) with supplementary condition (2.20) and the outer equation (2.21) will be solved simultaneously, as a system, and the solution $(\Psi,\phi_{\rm out})$ will depend analytically on $\omega$. The reduced equation (2.19) will determine $\omega$ by assigning the remaining $o(n^{-1/2})$ term $\widehat{\omega}=\widehat{\omega}(n)$ in the expansion (1.27) of $\omega$, see Lemma 2.12. To solve this equation, it will be necessary to examine precisely how $\omega$ enters the error term $K_{\rm err}$, see (2.87). This is analogous to how the Plemelj formula enters the calculation, for shear flows and vortices, of the one-sided derivative $c^{\prime}(m_{0})$ of the wavespeed $c$ with respect to the wavenumber $m$ as $\operatorname{Im}c\to 0^{+}$ at a limiting neutral mode $(\phi_{0},m_{0},c_{0})$. [^6] A limiting neutral mode is a non-trivial solution $(\phi_{0},m_{0},c_{0})$ of the Rayleigh equation (1.5) with $\operatorname{Im}c_{0}=0$ and which is, furthermore, approximable by solutions $(\phi,m,c)$ with $\operatorname{Im}c>0$. In [37], the notation $\alpha$ is used instead of $m$. In [52, 53] and [4, Chapter 4], $z$ is used instead of $c$. Of course, our calculation is different, and our unstable modes are not perturbations of a neutral mode.
@@
$Y_{w}:=\{F\in H^{2}(\mathbb{R})\colon\|F\|_{Y_{w}}:=\|(|F^{\prime\prime}|+|\xi||F^{\prime}|+\xi^{2}|F|)\|_{L^{2}_{w}}<\infty\}\,,$ (2.23)
@@
with parameter $q_{0}$ taking any value in $(0,1)$. For convenience we fix as $q_{0}:=1/2$. Define the outer function space
@@
  • Section 2.2. We analyze the outer operator $\frac{\mathrm{d}^{2}}{\mathrm{d}r^{2}}-\widetilde{k}$ on $\mathbb{R}^{+}$.
@@ 2.1. The inner equation
$\Upsilon^{\prime\prime}-(K_{0}+K_{2}\xi^{2})\Upsilon=QF\,,\qquad\int\Upsilon\overline{W}\,\mathrm{d}\xi=0\,,$ (2.29)
@@
We emphasize that the right-hand side of the inner equation (2.16) involves $K_{\rm err}\Psi$, so, in order to use Lemma 2.2 to solve (2.16), one also needs to control $K_{\rm err}$. This is taken into account in the solvability lemma of the entire projected system (2.18)–(2.21) in Lemma 2.11, where $K_{\rm err}$ is controlled using (2.75).
@@
It is well known (see [12, Section 12]) that the $L^{2}(\mathbb{R})$ eigenfunctions of the operator $\frac{\mathrm{d}^{2}}{\mathrm{d}y^{2}}-y^{2}$ are, up to constant multiple, precisely the Hermite functions $G_{m}(y)=\mathrm{e}^{-y^{2}/2}H_{m}(y)$, with corresponding eigenvalues $-(2m-1)$, $m\geq 1$. When $\zeta\in(-\pi,\pi)$, $L_{\zeta}$ therefore has $L^{2}$ eigenfunctions $G_{m}(\mathrm{e}^{i\zeta/2}x)$ with corresponding eigenvalue $-(2m-1)\mathrm{e}^{i\zeta}$, $m\geq 1$. In Lemma 2.5, we verify that these eigenvalues, produced by Wick rotation, are algebraically simple and exhaust the spectrum.
@@
We first note that the equation implies that $\phi\in H^{2}_{\rm loc}(\mathbb{R})$. We multiply the equation by $e^{-i\zeta}\mu\overline{\phi}$, where $\mu:\mathbb{R}\to[0,\infty)$ is compactly supported and Lipschitz, and integrate by parts:
$(\operatorname{Re}\mathrm{e}^{-i\zeta})\int\mu|\phi^{\prime}|^{2}+\operatorname{Re}(\mathrm{e}^{-i\zeta}\int\mu^{\prime}\phi^{\prime}\overline{\phi})+(\operatorname{Re}\mathrm{e}^{i\zeta})\int\mu x^{2}|\phi|^{2}$ $=\operatorname{Re}\left(\mathrm{e}^{-i\zeta}\int\mu f\overline{\phi}-\lambda\mu|\phi|^{2}\right)\,.$ (2.39)
@@
$\chi(x):=\begin{cases}1\qquad&|x|\leq 1,\\ 2-|x|&x\in(1,2),\\ 0&|x|\geq 2\,,\end{cases}$
so that $|\mu^{\prime}|^{2}\leq 4\mu/R^{2}$. We split the boundary term, involving $\mu^{\prime}\phi^{\prime}\overline{\phi}=\mu^{\prime}\mu^{-1/2}\phi^{\prime}\times\mu^{1/2}\overline{\phi}$, using Young’s inequality. We take $R\to\infty$ to obtain
$\int|\phi^{\prime}|^{2}+x^{2}|\phi|^{2}\lesssim(\operatorname{Re}\mathrm{e}^{i\zeta})^{-1}\int\left(|f|^{2}+|\lambda||\phi|^{2}\right).$ (2.41)
@@
Next, we multiply the equation by $\mathrm{e}^{-i\zeta}x^{2}\mu\overline{\phi}$ and integrate by parts to obtain
@@
$\int x^{2}|\phi^{\prime}|^{2}+\int x^{4}|\phi|^{2}\lesssim(\operatorname{Re}\mathrm{e}^{i\zeta})^{-2}\int|f|^{2}+|\lambda|^{2}|\phi|^{2}\,,$ (2.43)
@@
[…]
As a consequence of Lemma 2.4, the operator $L_{\zeta}$ has compact resolvent, i.e., there exists $\lambda$ in the resolvent set $\varrho(L_{\zeta})$ (which is non-empty, as it includes $\{\operatorname{Re}\lambda>0\}$) such that $R(\lambda,L_{\zeta}):L^{2}\to L^{2}$ is compact. Therefore, $\sigma(L_{\zeta})$ consists of isolated eigenvalues with finite multiplicity, see [30, p. 187, Theorem 6.29]. Thus, to classify the spectrum, it will suffice to classify eigenvalues.
@@
Let $\zeta\in(-\pi/2,\pi/2)$. Then $\sigma(L_{\zeta})$ consists of algebraically simply eigenvalues $-\mathrm{e}^{i\zeta}(2m-1)$, $m\geq 1$, with corresponding eigenspaces spanned by $G_{m}(\mathrm{e}^{i\zeta}\cdot)$.
@@
First, we prove algebraic simplicity by continuity. To be precise, we note that the map $\zeta\mapsto L_{\zeta}:(-\pi/2,\pi/2)\to\mathcal{B}(D(L_{0});L^{2})$ is continuous, and we fix any eigenvalue of $L_{0}$, namely, $\lambda_{m}=-(2m-1)$, $m\geq 1$, which is known to be algebraically simple because its geometric multiplicity is one and $L_{0}$ is self-adjoint. The curve $\zeta\mapsto\mathrm{e}^{i\zeta}\lambda_{m}$ is a curve of eigenvalues of $L_{\zeta}$, with each of the eigenvalues isolated in $\sigma(L_{\zeta})$. Then [30, p. 212-213, Theorem 3.16] yields that the algebraic multiplicity of $\mathrm{e}^{i\zeta}\lambda_{m}$ is constant along the curve.
@@
where $\phi(x)=w(y)$, $y=\sqrt{2}x$, and $\lambda=2\mathrm{e}^{i\zeta}a$. The solution space is the linear span of the parabolic cylinder functions $U(a,z)$, $V(a,z)$ evaluated on the line $z=\mathrm{e}^{i\zeta/2}y$. We follow the conventions of [12, Section 12.2], in which $U$ and $V$ are linearly independent entire solutions to $U^{\prime\prime}(z)-z^{2}U(z)/4=aU(z)$. We use the asymptotic expansions for $|z|\to\infty$ in certain sectors:
@@
where $\Gamma$ denotes the Gamma function, and $\delta>0$ is arbitrarily small (see [12, Section 12.9] and the comprehensive discussion in [50] for details). Hence, $V$ blows up as $|z|\to\infty$ along the ray ${\rm Arg}\,z=\zeta/2$ (that is, as $x\to\infty$). Meanwhile, $U$ blows up as $|z|\to\infty$ along the ray ${\rm Arg}\,z=\zeta/2+\pi$ (that is, as $x\to-\infty$) only when $a\neq-(m+1/2)$, $m\geq 0$, so that $1/\Gamma(a+1/2)\neq 0$. When $a=-(m+1/2)$, $m\geq 0$, $1/\Gamma(a+1/2)$ has a simple zero which cancels the blowing-up part in the asymptotic expansion for $U(a,z)$. In that setting, $U(a,z)$ is a constant multiple of a Hermite function $G_{m+1}(z/\sqrt{2})$ (see [12, Eq. 12.7.2]). ∎
To complete the proof of Lemma 2.2, we will appeal to the Fredholm theory. For $\lambda\in\sigma(L_{\zeta})$, a theorem concerning isolated eigenvalues of finite multiplicity (see [30, p. 239, Theorem 5.28]) guarantees that $\lambda-L_{\zeta}$ is Fredholm. In particular, its range is closed. We recall the following.
@@
(Banach closed range theorem [55, Section VII.5]) Let $L\colon D(L)\subset\mathcal{H}\to\mathcal{H}$ denote a closed linear operator on some Hilbert space $\mathcal{H}$ such that ${\rm range}\;L$ is closed, and let $L^{*}$ denote its adjoint operator, defined on $D(T^{*})\subset\mathcal{H}$. Then
${\rm range}\;L=(\ker L^{*})^{\perp}\qquad\text{ and }\qquad{\rm range}\;L^{*}=(\ker L)^{\perp}.$
@@
Let $\zeta\in(-\pi/2,\pi/2)$ and $\lambda=-e^{i\zeta}(2m-1)$, $m\geq 1$, be an eigenvalue of $L_{\zeta}$. The adjoint (see [30, p. 167-168]) of $\lambda-L_{\zeta}$ is
$(\lambda-L_{\zeta})^{*}=\overline{\lambda}-L_{-\zeta}\,.$ (2.50)
@@
$\mathrm{range}\,(\lambda-L_{\zeta})$ $=(\mathrm{ker}\,(\overline{\lambda}-L_{-\zeta}))^{\perp}$ (2.51)
@@
=$=\left\{F\in L^{2}\colon\int F(y)G_{m}(\mathrm{e}^{i\zeta/2}y)\,\mathrm{d}y=0\right\}=:H\,.$
@@
$V:=\left\{\Psi\in D(L_{\zeta})\colon\int_{\mathbb{R}}\Psi(y)\overline{G_{m}(\mathrm{e}^{i\zeta}y}\,\mathrm{d}y=0\right\}\,,$ (2.52)
@@
$\int G_{m}^{2}(\mathrm{e}^{i\zeta/2}y)\,\mathrm{d}y\neq 0\,.$ (2.53)
@@
Suppose $\int G_{m}^{2}(\mathrm{e}^{i\zeta/2}y)\,\mathrm{d}y=0$. By the characterization of $H$ in (2.51), $G_{m}(\mathrm{e}^{i\zeta/2}\cdot)$ belongs to the range of $\lambda-L_{\zeta}$, $\lambda=-\mathrm{e}^{i\zeta}(2m-1)$. Hence, there exists $\Psi\in D(L_{\zeta})$ satisfying $(\lambda-L_{\zeta})^{2}\Psi=(\lambda-L_{\zeta})G_{m}(\mathrm{e}^{i\zeta}\cdot)=0$. In particular, either $\Psi$ is a constant multiple of $G_{m}(\mathrm{e}^{i\zeta}\cdot)$, which is a contradiction, or otherwise the eigenvalue $\lambda$ has a non-trivial Jordan block with generalized eigenfunction $\Psi$, which is also a contradiction. ∎

### v3

@@ 1.3. Nonlinear instability
coupled with the one-dimensional reduced equation (IQ)G=0(I-Q)G=0, that is,
$\int GW=0\quad\Leftrightarrow\quad\int\left(\widetilde{X}_{\rm in}K_{\rm err}(W+\Psi)-R\phi_{\rm out}X_{\rm out}^{\prime\prime}-2n^{-3/4}R\phi_{\rm out}^{\prime}X_{\rm out}^{\prime}\right)W\,\mathrm{d}\xi=0\,.$ (2.19)
@@
$\int\Psi W^{*}=0\,$, (2.20)
for example, and when $K_{\rm err}$ is small. The latter condition dictates some constraints on $\ell_{\rm in},\ell_{\rm out}$ (in addition to (2.2)), see the discussion below Lemma 2.11 for details.
Thus far, we have neglected the outer equation (recall (2.4)). It will be extended to $\mathbb{R}^{+}$:
@@
and, after we analyze $k$ in the region $|r-r_{0}|\gtrsim n^{-3/4}$, where $\widetilde{k}$ is defined in (2.55) and denotes an appropriate extension of $k$ guaranteeing coercivity.
To complete the construction, the projected equation (2.18) with supplementary condition (2.20) and the outer equation (2.21) will be solved simultaneously, as a system, and the solution $(\Psi,\phi_{\rm out})$ will be holomorphic with respect to $\omega$. The reduced equation (2.19) will determine $\omega$ by specifying the remaining $o(n^{-1/2})$ term $\widehat{\omega}=\widehat{\omega}(n)$ in the expansion (1.28) of $\omega$, see Lemma 2.12. To solve this equation, it will be necessary to examine precisely how $\omega$ enters the error term $K_{\rm err}$, see (2.88). This is analogous to how the Plemelj formula enters the calculation, for shear flows and vortices, of the one-sided derivative $c^{\prime}(m_{0})$ of the wavespeed $c$ with respect to the wavenumber $m$ as $\operatorname{Im}c\to 0^{+}$ at a limiting neutral mode $(\phi_{0},m_{0},c_{0})$. [^6] A limiting neutral mode is a non-trivial solution $(\phi_{0},m_{0},c_{0})$ of the Rayleigh equation (1.5) with $\operatorname{Im}c_{0}=0$ and which is, furthermore, approximable by solutions $(\phi,m,c)$ with $\operatorname{Im}c>0$. In [33, 39], the notation $\alpha$ is used instead of $m$. In [54, 55] and [4, Chapter 4], $z$ is used instead of $c$. Of course, our calculation is different, and our unstable modes are not perturbations of a neutral mode.
@@
$Y_{w}:=\{F\in H^{2}(\mathbb{R})\colon\|F\|_{Y_{w}}:=\||F^{\prime\prime}|+|\xi||F^{\prime}|+\xi^{2}|F|\|_{L^{2}_{w}}<\infty\}\,$, (2.23)
@@
with parameter $q_{0}$ taking any value in $(0,1)$. For convenience we fix $q_{0}:=1/2$. We define the outer function space
@@
  • Section $2.$ We analyze the outer operator $\frac{\mathrm{d}^{2}}{\mathrm{d}r^{2}}-\widetilde{k}$ on $\mathbb{R}^{+}$.
@@ 2.1. The inner equation
$\Upsilon^{\prime\prime}-(K_{0}+K_{2}\xi^{2})\Upsilon=QF\,,\qquad\int\Upsilon W^{*}\,\mathrm{d}\xi=0\,$, (2.29)
@@
We emphasize that the right-hand side of the inner equation (2.16) involves $K_{\rm err}\Psi$, and so, in order to use Lemma 2.2 to solve (2.16), one also needs to control $K_{\rm err}$. This is taken into account in the solvability lemma of the entire projected system (2.18)–(2.21) in Lemma 2.11, where $K_{\rm err}$ is controlled using (2.76).
@@
It is well known (see [12, Section 12]) that the $L^{2}(\mathbb{R})$ eigenfunctions of the operator $\frac{\mathrm{d}^{2}}{\mathrm{d}y^{2}}-y^{2}$ are, up to constant multiple, precisely the Hermite functions $G_{m}(y)=\mathrm{e}^{-y^{2}/2}H_{m}(y)$, with corresponding eigenvalues (2m1)-(2m-1), $m\geq 1$. When $\zeta\in(-\pi,\pi)$, $L_{\zeta}$ therefore has $L^{2}$ eigenfunctions $G_{m}(\mathrm{e}^{i\zeta/2}x)$ with corresponding eigenvalue $(2m-1)\mathrm{e}^{i\zeta}$, $m\geq 1$. In Lemma 2.5, we verify that these eigenvalues, produced by Wick rotation, are algebraically simple and exhaust the spectrum.
@@
We first note that the equation implies that $\phi\in H^{2}_{\rm loc}(\mathbb{R})$. We multiply the equation by $\mathrm{e}^{-i\zeta}\mu\phi^{*}$, where $\mu:\mathbb{R}\to[0,\infty)$ is compactly supported and Lipschitz, and integrate by parts:
$(\operatorname{Re}\mathrm{e}^{-i\zeta})\int\mu|\phi^{\prime}|^{2}+\operatorname{Re}\left(\mathrm{e}^{-i\zeta}\int\mu^{\prime}\phi^{\prime}\phi^{*}\right)+(\operatorname{Re}\mathrm{e}^{i\zeta})\int\mu x^{2}|\phi|^{2}$ $=\operatorname{Re}\left(\mathrm{e}^{-i\zeta}\int\mu f\phi^{*}-\lambda\mu|\phi|^{2}\right)\,.$ (2.39)
@@
[…]
so that $|\mu^{\prime}|^{2}\leq 4\mu/R^{2}$. We split the boundary term, involving $\mu^{\prime}\phi^{\prime}\phi^{*}=\mu^{\prime}\mu^{-1/2}\phi^{\prime}\times\mu^{1/2}\phi^{*}$, using Young’s inequality. We take $R\to\infty$ to obtain
$\int\left(|\phi^{\prime}|^{2}+x^{2}|\phi|^{2}\right)\lesssim(\operatorname{Re}\mathrm{e}^{i\zeta})^{-1}\int\left(|f|^{2}+|\lambda||\phi|^{2}\right).$ (2.41)
@@
Next, we multiply the equation by $\mathrm{e}^{-i\zeta}x^{2}\mu\phi^{*}$ and integrate by parts to obtain
@@
$\int x^{2}|\phi^{\prime}|^{2}+\int x^{4}|\phi|^{2}\lesssim(\operatorname{Re}\mathrm{e}^{i\zeta})^{-2}\int\left(|f|^{2}+|\lambda|^{2}|\phi|^{2}\right),$ (2.43)
@@
The proof can be obtained by a diagonal argument using a sequence of increasing radii $R_{n}\to\infty$, as well as the compact embedding $H^{2}(-R_{n},R_{n})\subset L^{2}(-R_{n},R_{n})$ for each $n$, and the decay at $\pm\infty$ provided by $\|\cdot\|_{D(L_{\zeta})}$. ∎
As a consequence of Lemma 2.4, the operator $L_{\zeta}$ has compact resolvent, i.e., there exists $\lambda$ in the resolvent set $\varrho(L_{\zeta})$ (which is non-empty, as it includes $\{\operatorname{Re}\lambda>0\}$) such that $R(\lambda,L_{\zeta}):L^{2}\to L^{2}$ is compact. Therefore, $\sigma(L_{\zeta})$ consists of isolated eigenvalues with finite multiplicity, see [31, p. 187, Theorem 6.29]. Thus, to classify the spectrum, it will suffice to classify eigenvalues.
@@
Let $\zeta\in(-\pi/2,\pi/2)$. Then $\sigma(L_{\zeta})$ consists of algebraically simply eigenvalues $\mathrm{e}^{i\zeta}(2m-1)$, $m\geq 1$, with corresponding eigenspaces spanned by $G_{m}(\mathrm{e}^{i\zeta}\cdot)$.
@@
First, we prove algebraic simplicity by continuity. To be precise, we note that the map $\zeta\mapsto L_{\zeta}:(-\pi/2,\pi/2)\to\mathcal{B}(D(L_{0});L^{2})$ is continuous, and we fix any eigenvalue of $L_{0}$, namely, $\lambda_{m}=-(2m-1)$, $m\geq 1$, which is known to be algebraically simple because its geometric multiplicity is one and $L_{0}$ is self-adjoint. The curve $\zeta\mapsto\mathrm{e}^{i\zeta}\lambda_{m}$ is a curve of eigenvalues of $L_{\zeta}$, with each of the eigenvalues isolated in $\sigma(L_{\zeta})$. Then [31, p. 212-213, Theorem 3.16] yields that the algebraic multiplicity of $\mathrm{e}^{i\zeta}\lambda_{m}$ is constant along the curve.
@@
for $a\in\mathbb{C}$, where $\phi(x)=w(y)$, $y=\sqrt{2}x$, and $\lambda=2\mathrm{e}^{i\zeta}a$. The solution space is the linear span of the parabolic cylinder functions $U(a,z)$, $V(a,z)$ evaluated on the line $z=\mathrm{e}^{i\zeta/2}y$. We follow the conventions of [12, Section 12.2], in which $U$ and $V$ are linearly independent entire solutions to $U^{\prime\prime}(z)-z^{2}U(z)/4=aU(z)$. We use the asymptotic expansions for $|z|\to\infty$ in certain sectors:
@@
where $\Gamma$ denotes the Gamma function, and $\delta>0$ is arbitrarily small (see [12, Section 12.9] and the comprehensive discussion in [52] for details). Hence, $V$ blows up as $|z|\to\infty$ along the ray ${\rm Arg}\,z=\zeta/2$ (that is, as $x\to\infty$). Meanwhile, $U$ blows up as $|z|\to\infty$ along the ray ${\rm Arg}\,z=\zeta/2+\pi$ (that is, as $x\to$-$\infty$) only when $a\neq-(m+1/2)$, $m\geq 0$, so that $1/\Gamma(a+1/2)\neq 0$. When a=(m+1/2)a=-(m+1/2), $m\geq 0$, $1/\Gamma(a+1/2)$ has a simple zero which cancels the blowing-up part in the asymptotic expansion for $U(a,z)$. In that setting, $U(a,z)$ is a constant multiple of a Hermite function $G_{m+1}(z/\sqrt{2})$ (see [12, Eq. 12.7.2]). ∎
To complete the proof of Lemma 2.2, we will appeal to the Fredholm theory. For $\lambda\in\sigma(L_{\zeta})$, a theorem concerning isolated eigenvalues of finite multiplicity (see [31, p. 239, Theorem 5.28]) guarantees that $\lambda-L_{\zeta}$ is Fredholm. In particular, its range is closed. We recall the following.
@@
(Banach closed range theorem [57, Section VII.5]) Let $L\colon D(L)\subset\mathcal{H}\to\mathcal{H}$ denote a closed linear operator on some Hilbert space $\mathcal{H}$ such that ${\rm range}\;L$ is closed, and let $L^{*}$ denote its adjoint operator, defined on $D(L^{*})\subset\mathcal{H}$. Then
[…]
@@
Let $\zeta\in(-\pi/2,\pi/2)$ and $\lambda=-\mathrm{e}^{i\zeta}(2m-1)$, $m\geq 1$, be an eigenvalue of $L_{\zeta}$. The adjoint (see [31, p. 167-168]) of $\lambda-L_{\zeta}$ is
$(\lambda-L_{\zeta})^{*}=\lambda^{*}-L_{-\zeta}\,.$ (2.50)
@@
$\mathrm{range}\,(\lambda-L_{\zeta})$ $=(\mathrm{ker}\,(\lambda^{*}-L_{-\zeta}))^{\perp}$ (2.51)
@@
$=\left\{F\in L^{2}\colon\int F(y)G_{m}(\mathrm{e}^{i\zeta/2}y)\,\mathrm{d}y=0\right\}=:H\,.$
@@
$V:=\left\{\Psi\in D(L_{\zeta})\colon\int_{\mathbb{R}}\Psi(y)(G_{m}(\mathrm{e}^{i\zeta}y))^{*}\,\mathrm{d}y=0\right\}\,$, (2.52)
@@
$\int G_{m}(\mathrm{e}^{i\zeta/2}y)^{2}\,\mathrm{d}y\neq 0\,.$ (2.53)
@@
We can calculate (2.53) by shifting the integration contour to the real axis:
$\mathrm{e}^{i\zeta/2}\int_{\mathbb{R}}G_{m}(\mathrm{e}^{i\zeta/2}y)^{2}\,\mathrm{d}y=\int_{\gamma}G_{m}(z)^{2}\,\mathrm{d}z=\int_{-\infty}^{+\infty}G_{m}(z)^{2}\,\,\mathrm{d}z>0\,$, (2.54)
where $\gamma$ is the contour parameterized by $\gamma(y)=\mathrm{e}^{i\zeta/2}y$, $y\in\mathbb{R}$. ∎